OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 20 square feet. Estimate the radius of the pool, to the nearest foot. A.14feet B.8feet C.11feet D.10feet

OpenStudy (mertsj):

Remember the area formula: A = pi(r^2)

OpenStudy (callisto):

let x be the radius. (pi) (x^2) = 20 x^2 = 6.366197724 x = 2.52 feet or x=-2.52 feet (rejected) @@.. not from the choice you gave us